when a mixture of acetone and propanal is burnt, 4.48 liters of carbon dioxide are produced at STP. What is the mass of water produced at the same time?
1
Expert's answer
2019-05-10T08:05:28-0400

Acetone and propanal are functional isomers having the same molecular formula C3H6O but different functional groups:



This means that both compounds have the same equation of combustion:


C3H6O + 4O2 = 3CO2 + 3H2O


Using ideal gas law pV=nRT we can find number of moles of carbon dioxide:


"n(CO_2) = \frac{pV}{RT} = \frac {101325Pa \cdot 0.00448m^3}{8.314\frac{J}{mol \cdot K}\cdot 273.15K} = 0.2 mol"


According to stoichiometry of the reaction


"n(H_2O) = n(CO_2) = 0.2 mol"


Thus mass of water produced in this reaction is


"m(H_2O) = n(H_2O)\cdot M(H_2O) = 0.2 mol \cdot 18\frac{g}{mol} = 3.6g"